# Fully Distributed Fog Load Balancing with Multi-Agent Reinforcement Learning

### v1

@@ VI Conclusion
This paper presents a MARL fully distributed load-balancing solution that optimizes end-to-end execution time by minimizing Fog nodes’ waiting delay. The agents in our design work independently, and can dynamically adapt to environment changes using transfer learning in a life-long learning process. Our findings demonstrate the effectiveness of the proposed distributed solution in minimizing waiting time while enhancing resource utilization fairness across heterogeneous fog resources. By leveraging distributed decision-making, MARL agents outperform centralized solutions and traditional baselines, leading to significant improvements in overall system performance. Furthermore, we evaluate a more realistic approach to observing the state of the environment through an interval-based multi-casting protocol. Considering the real cost of observing the environment has never been considered before, and instead, observations are assumed to be unrealistically available in real-time for every decision to make.
Unlike centralized load-balancing solutions, using fully distributed agents allows for a global-scale solution where groups of agents can work independently in small collaboration regions. For this reason, a primary focus for future research is to provide a real-world implementation to demonstrate the scalability and effectiveness of the proposed distributed load-balancing solution at a global scale. This can involve deploying independent agents in diverse Fog environments that span across distant geographical regions and heterogeneous network infrastructures. This research direction will bridge the gap between theoretical research and practical implementation, demonstrating the practical feasibility and effectiveness of the proposed solution at a global scale. In addition, testing with a real-world implementation can uncover potential limitations and constraints that can guide future research efforts in enhancing Fog load-balancing solutions.

### v2

@@ VI Conclusion
This paper presents a scalable Fog load balancing solution using fully distributed RL agents that independently learn to reduce workload accumulation in Fog nodes without exchanging learning or behavior information. Blindly optimizing the common objective of reducing Fog load, with independent strategies, allows our agents to reduce the average waiting delay, balance waiting times between Fog nodes, and provide fair resource utilization, leading to optimized end-to-end execution delay for IoT applications. Deploying an independent agent for each IoT AP leads to separate load distribution strategies, i.e., policies, that converge faster than learning a common policy for all APs. With limited training, our results show that learning independent policies outperforms explicitly learning a common policy by a single agent or CTDE-based distributed agents. Even with separate training, coordination between MARL agents implicitly leads to a common policy due to sharing network parameters, buffer experiences, or load distribution decisions.
Unlike centralized and coordination-based load-balancing solutions, fully distributed agents allow global-scale implementations. The independent agents can be grouped to manage smaller sets of Fog nodes from the global network, forming Fog regions that might overlap by shared Fog resources. The ability of these agents to optimize a common objective in an isolated region without coordination guarantees their ability to optimize that objective even when some Fog resources are shared with another region. Managing fewer Fog nodes in each region reduces the state and action spaces, which makes learning easier and faster. To reduce required training further, the agents use a lifelong learning framework to adapt faster to environmental changes while mitigating the computational cost of continuous training. To do this, the agents are fine-tuned using transfer learning as generation rates significantly increase in the environment. Transferring policy network weights and replay buffer experiences between sources and target learning tasks leads to faster convergences and better generalization than learning from scratch.
Each agent needs load information, i.e., for observations and rewards, from every Fog node in the network to make optimal load distribution decisions. Given communication latency, we can not assume the real-time availability of such information before every decision, especially with frequently generated workloads in large networks. Unfortunately, this assumption is common in the literature to report optimal performance that is not achievable in practice. Because the cost of collecting this information has never been considered before, we analyze a practical approach to observing Fog information in practical deployment. This analysis shows the trade-off between theoretical and practical performance by evaluating interval-based observations with realistic intervals from the Gossip multi-casting protocol. The findings demonstrate no impact on the single-agent solution and a slight deterioration in performance on the distributed agents. The unavoidable performance deterioration by independent agents is due to outdated observations, which make the environment non-stationary from the point of view of each agent, leading to harder learning.
For future work, adaptive observation intervals could be explored as an alternative to fixed intervals, where the intervals are adjusted based on demand fluctuations. This approach could reduce overhead and mitigate the impact of outdated observations. Moreover, observation intervals could vary for each agent based on its location in the network relative to the Fog nodes it manages. Additionally, a real-world implementation is essential to validate the scalability and effectiveness of the proposed fully distributed load-balancing solution in practice. Deploying independent agents across geographically dispersed Fog environments with heterogeneous resources would provide a more accurate and realistic evaluation, especially in public networks with rapidly changing demands. This research direction will help bridge the gap between theoretical results and practical performance, demonstrating the feasibility of the proposed approach in real-world scenarios.